OpenStudy (anonymous):

Use rational exponents to simplify. Write the answer in radical notation, if possible. ^12sqrta^10

OpenStudy (anonymous):

\[a^{\frac{10}{12}}\] \[a^{\frac{5}{6}}\]

jimthompson5910 (jim_thompson5910):

\[\sqrt[12]{a^{10}}=a^{\frac{10}{12}}=a^{\frac{5}{6}}=\sqrt[6]{a^5}\] So \[\sqrt[12]{a^{10}}=\sqrt[6]{a^5}\]